carmelle:

Which of the following statements is true only if triangles EFI and GFH are similar? A. 2FI = 3FH B. Lines EI and HG are parallel C. Points E, F, and G are collinear D. EI / FI = GH / FH

carmelle:

1 attachment
mhanifa:

What do you think?

applesunited:

dont know

carmelle:

@mhanifa wrote:
What do you think?
i dont know??

applesunited:

A maybe

mhanifa:

Can you test each option?

carmelle:

how do i do that?

applesunited:

uh yea how do you do that

mhanifa:

Start with easy ones, B and C.

applesunited:

i guess cross those out

mhanifa:

B. Lines EI and HG are parallel Are they parallel?

carmelle:

no

applesunited:

no

mhanifa:

C. Points E, F, and G are collinear Are they on same line?

carmelle:

no

applesunited:

no

mhanifa:

Both B and C are correct. The segments are parallel and the 3 points are collinear... But it doesn't come from the similarity of the triangles.

carmelle:

wait wym? theres only one correct option...

mhanifa:

What property do you know about similar triangles that can help to chose the only correct answer?

carmelle:

you mean, AAA, SAS, and SSS?

mhanifa:

The ratio of corresponding sides is same.

mhanifa:

You need to chose between A and D based on this property

carmelle:

ohh

carmelle:

since you said ratios, its D, right?

mhanifa:

Can you confirm if this is the ratio of corresponding sides?

carmelle:

no? i dont know how to do that..

mhanifa:

Can you match the corresponding sides of the two triangles?

mhanifa:

For example, EI and HG are corresponding sides

carmelle:

i dont understand.. sorry, can you explain it little bit more?

mhanifa:

The two triangles are similar, it means one of them is smaller copy of the other. Let one has sides 3, 4 and 5 and the other has sides 6, 8 and 10. Then 3 and 6 are corresponding as well as 4 and 8 or 5 and 10.

mhanifa:

You get same ratio when divide the matching sides: 6/3 = 2 8/4 = 2 10/5 = 2

mhanifa:

This is the ratio of corresponding sides

carmelle:

ohhh, i understand now, tysm

mhanifa:

Great. Now, can you list the sides of the bigger triangle?

carmelle:

do you mean this...? 3/6 = 1/2 4/8 = 1/2 10/5 = 1/2

mhanifa:

No, the one in the picture

carmelle:

what?? im confused again :')

lifewmylilweeb:

same

mhanifa:

EFI is the bigger triangle And its sides are: EF, FI and EI, right?

carmelle:

ohhh, you wanted me to name the sides... im dumb 🤦 alright... EF, FI and EI FH, FG, and HG

mhanifa:

No problem, its fine. I should have mentioned... Fine, can you tell which sides are matching between bigger and smaller triangles?

mhanifa:

I already mentioned one above

carmelle:

EI and HG EF and FH or FG?

mhanifa:

EF and FG. And the last pair is?

carmelle:

FI and FH

mhanifa:

If you divide each pair, you should get same number. Can you show this?

carmelle:

yea, thats what we just did, right?

mhanifa:

If the ratio is k, then it is: EI /HG = k EF /FG = k FI /FH = k

mhanifa:

Now ignore the middle one and make the others same

mhanifa:

EI/HG = FI/FH

mhanifa:

Can this be converted into the option D?

carmelle:

yea, i think so

mhanifa:

So you have the answer now

carmelle:

thank youuuu<3

mhanifa:

No problem. Have a good time!